OpenStudy (anonymous):

subtract and simplify (4-i)-(5+2i)=

OpenStudy (anonymous):

type answer in a+bi

OpenStudy (hba):

I think you can subtract.

OpenStudy (n00bstyle):

It's just a 'simple' subtraction, you can ignore the brackets: 4-i-5-2i

OpenStudy (anonymous):

(4-i)-(5+2i)=4-i-5-2i then 4-5-i-2i=-1-3i

OpenStudy (anonymous):

answer is -1-3i

OpenStudy (n00bstyle):

When it comes down to multiplying or dividing, you have to apply other rules !

OpenStudy (anonymous):

thank you one more

OpenStudy (n00bstyle):

https://www.youtube.com/watch?v=KPSj4-76eEc

OpenStudy (anonymous):

thnkx 2